OpenStudy (anonymous):

HELP ME WITH THIS 1) |3p|=3p 2) |7y-3|=3-7y

hartnn (hartnn):

do you know if \(|a|=b \\ \implies a=b \quad or \quad a=-b\) ?

hartnn (hartnn):

so if i have \(|a|=a\) i will have , a=a or a=-a a=-a will give me a= 0 same for the case of |a| =-a then a=-a or a=a , so, a=0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

A method is to square both sides to remove the modulus. (|3p|)^2 = (3p)^2 Another method is like what hartnn did;)

hartnn (hartnn):

@Farheen28 are you getting these methods ? ask doubts...

OpenStudy (anonymous):

so the answer is 0 ?

hartnn (hartnn):

for both, YES

hartnn (hartnn):

i mean 3p=0 gives p=0 but 7y-3 =0 gives y= ... ?

OpenStudy (anonymous):

3/7 right?

hartnn (hartnn):

correct! :)

OpenStudy (anonymous):

won't we solve it by separating the cases like 7y-3=3-7y and another would be -7y+3=3-7y ? @hartnn

hartnn (hartnn):

we have to solve like that only thats why i gave you example with "a" -7y+3=3-7y is always true and will give you nothing 7y-3=3-7y will give you y=3/7 :)

OpenStudy (anonymous):

oh i got it. but for the first one the answer in the book is p>=0, how is this?

hartnn (hartnn):

thats because of the definition of |...| |negative stuff| = +stuff (absolute sign makes the quantity positive) so, say if your p is negative, like -1 |-3| = -3 which is NOT TRUE , because |-3| actually = 3 hence there is a restriction on p that it must be positive or 0 hence p>=0 got it ?

OpenStudy (anonymous):

yes :D thanks a ton

hartnn (hartnn):

welcome ^_^